# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ D.3.1. Critical point in $I_{\alpha}^{\prime}$
For convenience, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. By (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b), $\Delta H(r,\check{t})=O(\check{t}^{-3/2})$, $\check{t}\to+\infty$. The statement $\Delta G=O(1/|\check{t}|)$, $\check{t}\to-\infty$, follows immediately from (D.1). Next, suppose $\check{t}\to+\infty$. Then
@@
(E.1) $\begin{split}&\Delta G(r,\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s.\end{split}$
@@
In view of (D.1), the integral in (E.1) is split into three
@@
(E.2) $\begin{split}&\bigg{(}\int_{-c}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s\\ &=J_{1}+J_{2}+J_{3}.\end{split}$
@@
Breaking up the integral over $[-c,\check{t}-1]$ into the integrals over $[-c,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{1}=O(1/\check{t})$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{3}=O(1/\check{t})$.
Also, from (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) we get
(E.3) $|H_{\check{t}}^{\prime}(\check{t}+r)-H_{\check{t}}^{\prime}(\check{t})|\leq O(\check{t}^{-3/2})+c\big{|}\partial_{\check{t}}\big{[}\chi(\epsilon(\check{t}+\check{q}))\check{t}^{-1/2}\big{]}\big{|}=O(\check{t}^{-3/2}).$
Therefore, using $|s-\check{t}|\leq 1$ gives
(E.4) $\partial_{s}\big{[}\chi_{1\text{D}}(\epsilon(s-\check{t}))\Delta H(r,s)\big{]}=\epsilon O(\check{t}^{-3/2})+O(\check{t}^{-3/2}),\ \check{t}\to+\infty.$
Combining the three estimates proves that the integral in (E.1) is $O(1/\check{t})$. Combining the results finishes the proof of the estimate of $\Delta G$.
The derivative $\partial_{r}\Delta G(r,\check{t})$ in (6.16) is estimated in exactly the same fashion. In particular, we use $\partial_{r}\Delta H(r,\check{t})=\partial_{r}H(r+\check{t})=O(\check{t}^{-3/2})$ as $\check{t}\to+\infty$.
To prove (7.6), use (E.1) to write
@@
(E.5) $\begin{split}&\Delta G(r,\check{t}+h)-\Delta G(r,\check{t})\\ &=\bigg{(}\int_{-1}^{1}+\int_{|s|>1}\bigg{)}\bigg{[}\frac{\chi_{1\text{D}}(\epsilon(s-h))}{s-h}-\frac{\chi_{1\text{D}}(\epsilon s)}{s}\bigg{]}\Delta H(r,s+\check{t})\text{d}s\\ &=:J_{1}(\check{t},h)+J_{2}(\check{t},h).\end{split}$
When $0<\epsilon\ll 1$ and $|s|<1$, we have $\chi_{1\text{D}}(\epsilon(s-h))\equiv\chi_{1\text{D}}(\epsilon s)\equiv 1$. Then
(E.6) $\begin{split}J_{1}(\check{t},h)&=\bigg{[}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{]}\bigg{|}_{s=-1}^{1}-\int_{-1}^{1}\Delta H_{\check{t}}^{\prime}(r,s+\check{t})\ln\bigg{|}\frac{s-h}{s}\bigg{|}\text{d}s\\ &=:J_{11}(\check{t},h)+J_{12}(\check{t},h).\end{split}$
By (LABEL:H_props_2_a) with $k=0$ and (LABEL:H_props_2_b),
(E.7) $\begin{split}\bigg{|}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{|}\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}},\ s=\pm 1.\end{split}$
To bound all the remaining terms, we can look at them individually instead of estimating differences at $\check{t}+r$ and $\check{t}$ as above. Indeed,
(E.8) $\begin{split}|J_{12}(\check{t},h)|&\leq\int_{-1}^{1}\big{|}H^{\prime}(s+\check{t})\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\int_{-1}^{1}(1+|s+\check{t}|)^{-3/2}\big{|}\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\frac{h}{(1+|\check{t}|)^{3/2}}\int_{-1/h}^{1/h}\big{|}\ln|1-(1/u)|\big{|}\text{d}u\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}}\end{split}$
and
(E.9) $\begin{split}&|J_{2}(\check{t},h)|\leq ch\int_{|s|>1}\bigg{(}\frac{1}{s^{2}}+\frac{\epsilon}{|s|}\bigg{)}\frac{1}{(1+|s+\check{t}|)^{3/2}}\text{d}s\\ &\leq ch(|\check{t}|^{-3/2}+\epsilon|\check{t}|^{-1}).\end{split}$
Combining the results finishes the proof.
@@
Pick any $A_{1}>0$. Suppose $|\check{t}|<A_{1}$, $|\epsilon\check{q}-P(0)|<\epsilon^{1/2}A_{1}$, and $|\theta|<\epsilon^{1/2}A_{1}$. Under the assumptions of Theorem 3.6 one has
(E.10) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=\mu G_{0}(\check{t},\check{q})+O(\epsilon^{1/2}).\end{split}$
Partition the interval $[-A,A]$ into the union of $K\gg 1$ non-overlapping intervals of length $2A/K$. Let these intervals be denoted $B_{k}$, and $s_{k}$ be their centers, $k=1,2,\dots,K$. Recall that $P(\alpha)$ is associated with $x_{0}\in\mathcal{S}$. Clearly,
@@
Here we have used that $A>0$ is fixed and $|s_{k}|<A$. Otherwise, the term $O(1/K)$ may grow with $A$. The same applies to many other bounds and big-$O$ terms in this section. Using that $H_{0}$ is compactly supported and $|\check{t}|<A$, we get from (7.16)
(E.12) $\begin{split}&G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})=-\int_{-c}^{c}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\partial_{\check{t}}H_{0}(w+\check{t},\check{q})\text{d}w,\ |\check{t}|<A.\end{split}$
Hence, by assumption 3.4(1) and (7.16), $|\partial_{\check{t}}H_{0}(w+\check{t},\check{q})|<c$ and
(E.13) $\begin{split}&|G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})|<c\int_{-c}^{c}\bigg{|}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\bigg{|}\text{d}w<ca\ln(1/a),\\ &0<a<1,\ |\check{t}|<A.\end{split}$
Furthermore, assumption 3.4(1) implies
@@
Hence
@@
where $s=\epsilon^{1/2}\mu j_{1}\in B_{k}$ and $|\theta|<\epsilon^{1/2}A$. The equation (E.10) applies because by the assumptions of the lemma,
@@
There are $|B_{k}|/(\epsilon^{1/2}\mu)$ values of $j_{1}$ such that $\epsilon^{1/2}\mu j_{1}\in B_{k}$. Sum (E.18) with $s=\epsilon^{1/2}\mu j_{1}$ to obtain
(E.19) $\begin{split}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\theta,\epsilon\bigg{)}\\ =&\epsilon^{1/2}\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},u_{\epsilon}+P^{\prime}(0)\mu j_{1}\bigg{)}\\ &+|B_{k}|\big{[}O(1/K^{1/2})+O(\epsilon^{1/2})\big{]},\ u_{\epsilon}:=\frac{P(\epsilon^{1/2}s_{k})-P^{\prime}(0)\epsilon^{1/2}s_{k}}{\epsilon}.\end{split}$
It follows from (7.16) and (E.15) that $G_{0}$ is 1-periodic in its last argument and Hölder continuous.
From (4.8) and (D.30),
$P^{\prime}(\alpha)=[\Phi(x(\alpha),\alpha)]_{\alpha}^{\prime}=\Phi_{x}^{\prime}(x(\alpha),\alpha)x_{\alpha}^{\prime}(\alpha)+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)=\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).$
By the assumption of the theorem, $P^{\prime}(0)\mu$ is irrational, so the points $P^{\prime}(0)\mu j_{1}$ are uniformly distributed mod 1 [35, Exercise 5.6]. Take the limit $\epsilon\to 0$ in (E.19) and use [35, Corollary 1.2] to find
@@
Clearly the same arguments apply when we add $\Phi_{x}^{\prime}\check{x}$ to the first argument of $G$. Therefore, combining (E.21), (7.14) and (7.2), proves (7.15).
@@
Fix any $A_{1}$. Throughout the proof we assume $|\check{t}|<A_{1}$. Fix any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and denote
@@
In view of (6.11), (6.14), and (E.22), we estimate
(E.24) $\begin{split}&\bigg{|}\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)\big{[}a_{0}(\theta,\epsilon j_{2})-a_{0}(\theta,\epsilon\check{q})\big{]}(j_{2}-\check{q})_{+}^{1/2}\bigg{|}\\ &\leq c\epsilon\sum_{j_{2}=1}^{1/\epsilon}\frac{1}{j_{2}^{2}}j_{2}^{3/2}=O(\epsilon^{1/2}).\end{split}$
@@
(E.25) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(\theta,\epsilon\check{q})\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
Using the assumptions of Lemma E.1, we find
(E.26) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\big{[}a_{0}(0,P(0))+O(\epsilon^{1/2})\big{]}\\ &\times\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2})\\ =&a_{0}(0,P(0))\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
@@
(E.29) $\begin{split}\sum_{j_{2}}\mathcal{Q}_{c}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
Therefore, by (E.26) and (E.29),
(E.30) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(0,P(0))\int_{\mathbb{R}}\frac{1}{\check{w}}\sum_{j_{2}}h^{\prime}(\check{w}+\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}\text{d}\check{w}\\ &+O(\epsilon^{1/2}).\end{split}$
Further,
@@
where $s\in\text{supp}\,\varphi_{\alpha}$, which implies (see (6.10))
(E.32) $\begin{split}&\partial_{u}^{k}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}=O(\epsilon),\ k=0,1,\\ &|\partial_{u}^{2}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}|<c,\ u\in\mathbb{R},\theta=O(\epsilon^{1/2}),\end{split}$
Similarly to (E.22) and (E.23), pick any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and define
(E.33) $\begin{split}\mathcal{Q}_{\epsilon}(u):=\int_{\mathbb{R}}\frac{1}{\check{w}}\Delta\bar{h}_{\epsilon}^{\prime}(\check{w}+u)\text{d}\check{w},\ \Delta\bar{h}_{\epsilon}(u):=\bar{h}(u+\epsilon)-\bar{h}(u).\end{split}$

### v2

@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
For convenience, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. By (6.17) and (7.5),
@@
(E.1) $\begin{split}&\Delta G(r,\check{t})=G(r+\check{t})-G(\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s.\end{split}$
Begin with (7.7). Adding (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) yields $|\Delta H(r,\check{t})|\leq c|r|(1+|\check{t}|)^{-3/2}$. The bound $|\Delta G|\leq c|r|/|\check{t}|$, $\check{t}\to-\infty$, follows from (D.1).
@@
Suppose $\check{t}\to+\infty$. Using (D.1), the integral in (E.1) is split into three
@@
(E.2) $\begin{split}&\bigg{(}\int_{-c}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s\\ &=:J_{1}+J_{2}+J_{3}.\end{split}$
@@
Breaking up the integral over $[-c,\check{t}-1]$ into the integrals over $[-c,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $|J_{1}|\leq c|r|/\check{t}$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $|J_{3}|\leq c|r|\check{t}^{-3/2}\ln\check{t}$.
Using the function $(\chi_{1\text{D}}(\epsilon\check{w})-1)/\check{w}$, we obtain
@@
(E.3) $\begin{split}&J_{2}=J_{2}^{\prime}+O(\epsilon|r|\check{t}^{-3/2}),\\ &J_{2}^{\prime}=\int_{-1}^{1}\frac{1}{\check{w}}\Delta H(r,\check{w}+\check{t})\text{d}\check{w}=\int_{-1}^{1}\frac{1}{\check{w}}\int_{w}^{r+w}\partial_{\check{t}}H(s+\check{t})\text{d}s\text{d}\check{w}.\end{split}$
To estimate $J_{2}^{\prime}$, we may assume without loss of generality that $10<r<$ 1 and rewrite $J_{2}^{\prime}$ as follows
(E.4) $\begin{split}J_{2}^{\prime}=&\int_{-1}^{-1+r}\partial_{\check{t}}H(s+\check{t})\int_{-1}^{s}\frac{\text{d}\check{w}}{\check{w}}\text{d}s+\int_{-1+r}^{1}\partial_{\check{t}}H(s+\check{t})\int_{s-r}^{s}\frac{\text{d}\check{w}}{\check{w}}\text{d}s\\ &+\int_{1}^{1+r}\partial_{\check{t}}H(s+\check{t})\int_{s-r}^{1}\frac{\text{d}\check{w}}{\check{w}}\text{d}s.\end{split}$
Adding (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) yields $\partial_{\check{t}}H(\check{t})=O(\check{t}^{-3/2})$. After simple but tedious calculations, (E.4) implies:
(E.5) $\begin{split}|J_{2}^{\prime}|\leq&c\check{t}^{-3/2}\bigg{[}\int_{-1}^{-1+r}|\ln|s||\text{d}s+\int_{-1+r}^{1}\bigg{|}\ln\frac{|s|}{|s-r|}\bigg{|}\text{d}s+\int_{1}^{1+r}|\ln|s-r||\text{d}s\bigg{]}\\ =&O(r\ln(1/r)\check{t}^{-3/2}).\end{split}$
Therefore, $J_{2}=O(r\ln(1/r)\check{t}^{-3/2})$.
The above argument also proves that $\Delta G(r,\check{t})$ is bounded on compact sets. Combining the estimates of $J_{1},J_{2},J_{3}$ finishes the proof of (7.7).
It is easy to check that
(E.6) $\begin{split}&\Delta G(r+h,\check{t})-\Delta G(r,\check{t})=\Delta G(h,r+\check{t}),\\ &\Delta G(r,\check{t}+h)-\Delta G(r,\check{t})=\Delta G(h,r+\check{t})-\Delta G(h,\check{t}),\end{split}$
Since $|r|\leq A_{2}$, (7.8) and (7.9) follow from (7.7).
@@
Recall that the parameter $A\gg 1$ used in Lemma 7.2 was selected above the equation (7.13).
@@
Pick any $A_{1}\gg 1$. Suppose $|\check{t}|<A_{1}$, $|\epsilon\check{q}-P(0)|<\epsilon^{1/2}A_{1}$, and $|\theta|<\epsilon^{1/2}A_{1}$. Under the assumptions of Theorem 3.6 one has
(E.7) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=\mu G_{0}(\check{t},\check{q})+O(\epsilon^{1/2})\end{split}$
uniformly in $\check{t}$, $\check{q}$, and $\theta$ restricted to the indicated sets.
Partition the interval $[-A,A]$ into the union of $K\gg 1$ non-overlapping intervals of length $2A/K$. Let these intervals be denoted $B_{k}$, and $s_{k}$ be their centers, $k=1,2,\dots,K$. Recall that $P(\alpha)$ is associated with $x_{0}=0_{2}\in\mathcal{S}$. Clearly,
@@
Here we have used that $A>0$ is fixed and $|s_{k}|<A$. Otherwise, the term $O(1/K)$ may grow with $A$. The same applies to many other bounds and big-$O$ terms in this section.
Observe that $G_{0}(\check{t},\check{q})$ in (7.18) is obtained from $G(\check{t},\check{q};x,\theta,\epsilon)$ by setting $x=0_{2}$, $\theta=0$, $\epsilon=0$ and $a_{0}(\theta,\epsilon j_{2})\equiv a_{0}(y_{0})$ in (6.13), (6.14) and (6.17). Therefore, $G_{0}$ satisfies (7.7):
@@
To study the smoothness of $G_{0}$ with respect to the second argument, denote similarly to (7.18):
(E.10) $\begin{split}&\check{G}_{0}(\check{t},\check{q}_{1},\check{q}_{2}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q}_{1},\check{q}_{2})}{\check{w}}\text{d}\check{w},\\ &\check{H}_{0}(\check{t},\check{q}_{1},\check{q}_{2}):=a_{0}(y_{0})\sum_{j_{2}\in\mathbb{Z}}h_{0}^{\prime}(\check{t}+\check{q}_{1}-j_{2})(j_{2}-\check{q}_{2})_{+}^{1/2}.\end{split}$
Then, $G_{0}(\check{t},\check{q})=\check{G}_{0}(\check{t},\check{q},\check{q})$. Furthermore,
(E.11) $\begin{split}G_{0}(\check{t},\check{q}+r)-G_{0}(\check{t},\check{q})=&\big{[}G_{0}(\check{t},\check{q}+r)-G_{0}(\check{t}-r,\check{q}+r)\big{]}\\ &+\big{[}\check{G}_{0}(\check{t},\check{q},\check{q}+r)-\check{G}_{0}(\check{t},\check{q},\check{q})\big{]}.\end{split}$
The difference in the first brackets is bounded as in (E.9).
Clearly, $G_{0}(\check{t},\check{q}+1)\equiv G_{0}(\check{t},\check{q})$. Hence, we can assume $10\leq\check{q}<$ 1. Using that $h_{0}$ is compactly supported and $|\check{t}|<A$, we obtain
(E.12) $\begin{split}\int_{\mathbb{R}}\frac{h_{0}^{\prime}(\check{w}+\check{t}+\check{q}-j_{2})}{\check{w}}\text{d}\check{w}=O(j_{2}^{-2}),\ j_{2}\to\infty.\end{split}$
Also,
[…]
The estimates (E.12), (E.13) and (E.10) imply that the difference in the second brackets in (E.11) is $O(r^{1/2})$. Together with (E.9) this gives
@@
where $s=\epsilon^{1/2}\mu j_{1}\in B_{k}$ and $|\theta|<\epsilon^{1/2}A$. The equation (E.7) applies because the assumptions of Lemma 7.2 imply
@@
There are $|B_{k}|/(\epsilon^{1/2}\mu)$ values of $j_{1}$ such that $\epsilon^{1/2}\mu j_{1}\in B_{k}$. Substitute $s=\epsilon^{1/2}\mu j_{1}$ into (E.17) and sum with respect to $j_{1}$ to obtain
(E.18) $\begin{split}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\theta,\epsilon\bigg{)}\\ =&\epsilon^{1/2}\mu\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},u_{\epsilon}+P^{\prime}(0)\mu j_{1}\bigg{)}\\ &+|B_{k}|\big{[}O(1/K^{1/2})+O(\epsilon^{1/2})\big{]},\ u_{\epsilon}:=\frac{P(\epsilon^{1/2}s_{k})-P^{\prime}(0)\epsilon^{1/2}s_{k}}{\epsilon}.\end{split}$
The term $O(1/K^{1/2})$ is uniform with respect to $1\leq k\leq K$ and $0<\epsilon\ll 1$. The term $O(\epsilon^{1/2})$ is uniform with respect to $1\leq k\leq K$ and $K\geq 1$. It follows from (7.18) and (E.14) that $G_{0}$ is 1-periodic in its last argument and Hölder continuous.
From (4.8) and the definition $P(\alpha)=\Phi(x(\alpha),\alpha)$,
(E.19) $\begin{split}P^{\prime}(\alpha)&=[\Phi(x(\alpha),\alpha)]_{\alpha}^{\prime}=\text{d}_{x}\Phi(x(\alpha),\alpha)\cdot x_{\alpha}^{\prime}(\alpha)+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)\\ &=\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).\end{split}$
By the assumption of the theorem, $P^{\prime}(0)\mu$ is irrational, so the points $P^{\prime}(0)\mu j_{1}$ are well distributed mod 1 [38, Definition 5.1 and Example 5.2]. Take the limit $\epsilon\to 0$ in (E.18) and use [38, Theorem 5.1] to find
@@
Clearly the same arguments apply when we add $\text{d}_{x}\Phi\cdot\check{x}$ to the first argument of $G$, cf. (7.16). Therefore, combining (E.21), (7.16) and (7.4), proves (7.17).
@@
Let $A_{1}$, $\check{t}$, $\check{q}$ and $\theta$ be selected as in the statement of Lemma E.1. Fix any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and denote
@@
Substitute $\bar{h}(u)=h(u;\theta,\epsilon)$ of (6.13) into (E.22) to obtain $\mathcal{Q}(u;\theta,\epsilon)$. In view of (6.14), (6.17), and (E.22), we estimate
(E.24) $\begin{split}&\bigg{|}\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)\big{[}a_{0}(\theta,\epsilon j_{2})-a_{0}(\theta,\epsilon\check{q})\big{]}(j_{2}-\check{q})_{+}^{1/2}\bigg{|}\\ &\leq c\epsilon\sum_{j_{2}=1}^{1/\epsilon}\frac{1}{j_{2}^{2}}j_{2}^{3/2}=O(\epsilon^{1/2}).\end{split}$
@@
(E.25) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(\theta,\epsilon\check{q})\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
Since $|\check{t}|<A_{1}$, (E.23) implies that the sum in (E.25) is absolutely convergent and its value is uniformly bounded. Using the assumptions of Lemma E.1, we find
(E.26) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\big{[}a_{0}(0,P(0))+O(\epsilon^{1/2})\big{]}\\ &\times\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2})\\ =&a_{0}(0,P(0))\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
@@
(E.29) $\begin{split}\sum_{j_{2}\in J_{p}}\mathcal{Q}_{c}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
Due to the term $\epsilon/|u|$ in the bound in (E.28), it is essential that the sum in (E.29) is over a subset of the range $0<j_{2}-\check{q}<c/\epsilon$. Therefore, by (E.26) and (E.29),
(E.30) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(0,P(0))\int_{\mathbb{R}}\frac{1}{\check{w}}\sum_{j_{2}\in J_{p}}h^{\prime}(\check{w}+\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}\text{d}\check{w}\\ &+O(\epsilon^{1/2}).\end{split}$
It is obvious that if $\epsilon=0$ in (E.22) (i.e., if $\chi_{1\text{D}}$ is replaced by the constant function 1), then $\mathcal{Q}(u)=O(u^{-2})$, $u\in\mathbb{R}$. This implies that
(E.31) $\begin{split}\sum_{j_{2}-\check{q}>c/\epsilon}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
Hence, in (E.30), we can extend the summation to all of $j_{2}\in\mathbb{Z}$. Further,
@@
where $s\in\text{supp}\,\varphi_{\alpha}$. Here and in what follows, we drop the first argument of $K(x_{0}=0_{2},\alpha)$ from notation (see the paragraph following (7.3)). Using (6.13), introduce the variable
(E.33) $s_{1}=s_{1}(s;\theta,\epsilon)=\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon\mu\Phi^{\prime}(0)}.$
By the assumption in Theorem 3.6, $\mu\Phi_{\alpha}^{\prime}(0_{2},\alpha_{0})$ is irrational. Hence $\Phi^{\prime}(0)=\Phi_{\alpha}^{\prime}(0_{2},\alpha_{0})\not=0$. The assumption $|\theta|=O(\epsilon^{1/2})$ implies
(E.34) $s_{1}(s;\theta,\epsilon)-s=O(\epsilon^{1/2}),\ \partial s_{1}(s;\theta,\epsilon)/\partial s=1+O(\epsilon^{1/2}).$
Changing variable $s\to s_{1}$ in (6.13) gives
(E.35) $\begin{split}h(u;\theta,\epsilon)&=\mu\int_{\mathbb{R}}K(\theta+\epsilon\mu\zeta)\varphi_{\alpha}(\zeta)|\partial\zeta/\partial s|\varphi_{p}(\mu s\Phi^{\prime}(0)+u)\text{d}s,\ \zeta=\zeta(s;\theta,\epsilon).\end{split}$
To simplify the next step, in (E.35), the new variable $s_{1}$ is denoted $s$, and the old variable $s$ is denoted $\zeta$. Together with (7.18) this implies
(E.36) $\begin{split}&h(u;\theta,\epsilon)-\mu h_{0}(u)\\ &=\mu\int_{\mathbb{R}}\big{[}K(\theta+\epsilon\mu\zeta)\varphi_{\alpha}(\zeta)|\partial\zeta/\partial s|-K(0)\varphi_{\alpha}(s)\big{]}\varphi_{p}(\mu s\Phi^{\prime}(0)+u)\text{d}s.\end{split}$
Using that $\varphi_{\alpha},\varphi_{p}\in C_{0}^{2}(\mathbb{R})$ (see Assumption 3.4(1)) and (E.34), we conclude
(E.37) $\partial_{u}^{k}(h(u;\theta,\epsilon)-\mu h_{0}(u))=O(\epsilon^{1/2}),\ k=0,1,2.$
Similarly to (E.22) and (E.23), pick any $\bar{h}_{\epsilon}\in C_{0}^{2}(\mathbb{R})$ such that $\bar{h}_{\epsilon}^{(k)}=O(\epsilon^{1/2})$, $k=0,2$, and define
(E.38) $\begin{split}\mathcal{Q}_{\epsilon}(u):=\int_{\mathbb{R}}\frac{1}{\check{w}}\bar{h}_{\epsilon}^{\prime}(\check{w}+u)\text{d}\check{w}.\end{split}$